OpenStudy (anonymous):

Rewrite f(x) = 3(x − 2)2 + 1 from vertex form to standard form

OpenStudy (anonymous):

f(x) = 3x2 + 13 f(x) = 3x2−12x + 13 f(x) = 3x2 + 12x − 11 f(x) = 9x2− 36x + 37

OpenStudy (gabebae):

Standard form is Ax+Bx=C

Nnesha (nnesha):

that would be a straight line ^^ (vertical line )

Nnesha (nnesha):

\[\huge\rm f(x)=3(x-2)^2+1\] 2 is an exponent

OpenStudy (anonymous):

I was thinking it was f(x) = 3x2−12x + 13

OpenStudy (anonymous):

Is that right?

Nnesha (nnesha):

how did you get that ? post your work please

OpenStudy (anonymous):

I was guessing because it seems like a guessable answer but I actually have no idea

Nnesha (nnesha):

\[\huge\rm f(x)=3\color{Red}{(x-2)^2}+1\] (x-2)^2 is same as (x-2)(x-2) first we need to foil (x-2)^2

Nnesha (nnesha):

 multiply top terms by x